### v1

@@ 3.2 SN-driven destruction
where $M_{\rm g}$ is the local gas mass, $M_{{\rm d},i,j}$ is the local dust mass, $a_{0.1,j}$ is the dust grain size in units of $0.1\,\rm\mu m$, and $\delta_{{\rm SN},i}$ is the destruction efficiency. Small grains are more efficiently decelerated by drag forces, trapping them near the shock region where thermal sputtering can quickly destroy them (Nozawa et al. 2006). This functional form of the destruction efficiency has been constructed to capture this behaviour with grain size $a$. The mass of shocked gas is provided by the Sedov solution in a medium of homogeneous gas density, i.e. $M_{\rm 100}=6800E_{\rm SN,51}\,\rm M_{\odot}$, where $E_{\rm SN,51}$ is the energy of the SN explosion in units of $10^{51}\,\rm erg$. Multiple SNe are released at once over one time step $\Delta t$ (each individual SN engulfing the gas swept up by the previous explosion), following (see Hou et al. 2017):
@@
with $N_{\rm SN}$ the number of SNe. We stress that this destruction step by the SN blast solution does not explicitly destroy the returned ejecta from the stellar production, and destroys only the background dust material, hence, the dust condensation efficiency implicitly takes this term into account.
We use a different value of the dust destruction efficiency for carbonaceous and silicate grains with respectively $\delta_{\rm SN,C}=0.10$ and $\delta_{\rm SN,Sil}=0.15$ following the qualitative behaviour of Hu et al. (2019), where silicate grains are approximately destroyed 50 per cent more than carbonaceous grains.
@@ 3.3 Thermal sputtering
Dust is also destroyed by thermal sputtering, the ejection of atoms from grains by the transfer of kinetic energy from gas particles at temperatures high enough to overcome the energy barrier of the binding energy. We adopt the fitting form of Hu et al. (2019) to the thermal sputtering yields $Y_{\rm th}$ calculated by Nozawa et al. (2006), which relates to the thermal sputtering time scale as (recall that $t_{{\rm spu},i,j}=m_{{\rm gr},i,j}/|\dot{m}_{{\rm gr},i,j}|=a_{i,j}/(3|\dot{a}_{i,j}|)$ with $m_{{\rm gr},i,j}$ the mass of a single dust grain):
@@
Note that sputtering yields are approximately 3 times lower at $T\gtrsim 10^{6}\,\rm K$ for carbonaceous grains than for silicate grains (see also Draine & Salpeter 1979; Tielens et al. 1994). Hence, they differ from the widely adopted yields of Tsai & Mathews (1995) giving a unique sputtering time scale based on the intermediate values for silicates and carbonaceous grains of Draine & Salpeter (1979) and Tielens et al. (1994) (see Appendix C). Note that, using cosmological hydrodynamical simulations of galaxy clusters, Gjergo et al. (2018) and Vogelsberger et al. (2019) have shown that the dust mass content in observed galaxy clusters can be reproduced only for ten times lower sputtering yields than the canonical values. We defer this investigation to future work, but we immediately stress that thermal sputtering is already a sub-dominant destruction mechanism in our set of simulations compared to direct destruction by SNe. We have performed a G10LG simulation with sputtering time artificially increased by a factor of 10: there is no noticeable difference in any of the dust properties investigated in section 4.
@@ 3.4 Gas accretion
where $m_{{\rm gr},i,j}=s_{i}a_{j}^{3}4\pi/3$ is the grain mass, $s_{i}$ is the grain material density, $f_{X}$ is the mass fraction of the limiting element in the chemical composition of the grain, $u_{{\rm th},X}=(8k_{\rm B}T/(\pi m_{X}))^{1/2}$ is the gas thermal velocity, $m_{X}=A_{X}m_{\rm H}$ is the atomic mass, and $\rho_{X}$ is the total (gas plus dust phase) mass density of the limiting element. $\pi\bar{a}_{i,j}^{2}$ is the surface of the grain with Coulomb enhancement factor $E_{i}(a_{j})$ due to electrostatic effects caused by ionised gas interacting with charged grains (Weingartner & Draine 1999), which is integrated over the whole distribution of the grain sizes. In our case where the distribution is discretised over two bins, and assuming a Dirac distribution we simply get $\bar{a}_{i,j}=E_{i}(a_{j})a_{j}$. For carbonaceous grains, $X=i=\rm C$, however, for silicate grains, the limiting element depends on the grain composition, i.e.
@@
where $a_{{\rm 0.005},j}=a_{j}/5\,\rm nm$, $s_{3,i}=s_{i}/3\,\rm g\,cm^{-3}$, $\alpha$ is the sticking coefficient of gas particles onto dust, and $Z_{X}=\rho_{X}/\rho_{\rm g}$ is the mass fraction of the limiting element. Weingartner & Draine (1999) give values of $E(a)$ for a range of sizes of carbonaceous and silicate grains for three typical phases of the ISM. To mimic their complex dependencies, we simply adopt a value of $E=1$ for all kind of grains everywhere, except for large carbonaceous and small silicate grains in the cold neutral medium with $T<2\times 10^{4}\,\rm K$ and $n>10\,\rm H\,cm^{-3}$, for which we adopt $E_{\rm C}(0.1\,\mu\rm m)=0$, and $E_{\rm Sil}(5\,\rm nm)=10$.
@@
where $s=\log(n/\bar{n})$, $s_{0}=-\sigma_{s}^{2}/2$ and $\sigma_{s}=\log(1+b^{2}\mathcal{M}^{2})$, with $b$ the compression ratio (we take $b=0.4$ for mixed turbulence which is the same value used in the gravo-turbulent model for SF efficiency) and $\mathcal{M}$ the turbulent Mach number. The effective timescale $t_{\rm acc,eff}$ is obtained by integrating the mass accretion rate over the log-normal PDF for a given set of mean density, temperature and metallicity ($\bar{n},\bar{T},\bar{Z}$), hence:
@@
up to a maximum density $s_{\rm max}=\log(n_{\rm max}/\bar{n})$, where the maximum density is $n_{\rm max}=10^{4}\,\rm H\,cm^{-3}$ where grains starts to be significantly coated with water ice mantle ($n\geq 10^{3}\,\rm H\,cm^{-3}$, Cuppen & Herbst 2007; Hollenbach et al. 2009) suppressing the accretion of refractory material onto the grain surface. We use the value of the gas density and metallicity of a given cell for $\bar{n}$ and $\bar{Z}$, and we assume that the gas temperature is always $\bar{T}=100\,\rm K$ for gas that fulfills the conditions to trigger this unresolved log-normal PDF of density (see below) for simplicity. In addition, we assume the $T=\bar{T}$ and $Z=\bar{Z}$ for the sampled values of $n$ by the log-normal PDF. This leads to
@@
where $\rm Erfc$ is the complementary error function. This unresolved effective accretion timescale is estimated only for cells with gas density larger than $0.1\,\rm H\,cm^{-3}$, temperature lower than $10^{4}\,\rm K$, and Jeans length smaller than 4 times the local cell size, assuming a sticking coefficient of $\alpha_{\rm eff}=1/3$, which values are representative of the sticking coefficient obtained at temperatures of $T=100-1000\,\rm K$ (see Fig. 26). Otherwise the accretion time is that obtained for $\bar{n}$ (the local gas density) with the Le Bourlot et al. (2012) sticking coefficient (see Appendix D for a comparison of sticking coefficients).
@@
For carbonaceous grains, we further assume that carbon in the gas phase is fully molecular (CO) above a gas density of $n>10^{3}\,\rm H\,cm^{-3}$. Indeed, simulations (e.g. Safranek-Shrader et al. 2017; Clark et al. 2019; Hu et al. 2021; Katz et al. 2022) gives a transition to full CO gas at $n=2\times 10^{3}-10^{4}\,\rm H\,cm^{-3}$ depending on the CR ionisation rate. Note that while our value might seem fairly low (or corresponding to a very weak cosmic ray ionisation rate), our subgrid gas distribution neglects the fact that gas collapses further at smaller unresolved scales (higher density) and extends the log-normal PDF with a power-law tail (Vallini et al. 2018).
@@ 3.5 Shattering
Shattering is the process by which grains with sufficiently large velocities collide and fragment into smaller size grains. Therefore, it is a transfer of mass from the large to the small grain population that conserves the total dust mass. The corresponding timescale is obtained by considering the timescale for grain collisions $\tau_{\rm col}=(\sigma_{\rm D}\tilde{\sigma}n_{\rm D})^{-1}$, where $\tilde{\sigma}$ is the grain cross section, $\sigma_{\rm D}$ the dust grain velocity dispersion and $n_{\rm D}$ the dust number density, which boils down to (Aoyama et al. 2017):
@@
where $\mathcal{D}_{\rm L}$ and $\sigma_{\rm D,L}$ are, respectively, the dust-to-gas ratio and velocity dispersion of the large grains population. Typical velocities required to shatter grains are above a few $\rm km\,s^{-1}$ (Jones et al. 1996), which limit the effect of shattering to the diffuse phase of the ISM. Yan et al. (2004) provide grain-gas relative velocities for various phases of the ISM and show that large grains ($a_{\rm L}=0.1\,\rm\mu m$) reach the turbulent velocity of the forcing scale, i.e. $\sigma_{\rm D,L}\simeq 10\,\rm km\,s^{-1}$ in the warm ionised medium, $\simeq 1\,\rm km\,s^{-1}$ for the warm or cold neutral medium, and $\simeq 0.1-1\,\rm km\,s^{-1}$ for molecular clouds. We follow Granato et al. (2021) and adopt the following functional form:
@@
with $p_{\rm sh}=1$ for $n<1\,\rm H\,cm^{-3}$, and $p_{\rm sh}=1/3$ for $1<n<10^{3}\,\rm H\,cm^{-3}$, which is equivalent to having a dispersion velocity varying from $10\,\rm km\,s^{-1}$ to $1\,\rm km\,s^{-1}$ from the warm ionised phase ($1\,\rm H\,cm^{-3}$) to the molecular cloud density [^5] Note that considering the gas-drag timescale, and a Kolmogorov turbulence, the velocity dispersion of grains follow (Ormel et al. 2009; Hirashita & Aoyama 2019): $\sigma_{\rm gr}=1.2\mathcal{M}^{3/2}a_{0.1}^{1/2}s_{3}^{1/2}T_{4}^{1/4}n_{1}^{-1/4}\,\rm km\,s^{-1}$. Hence, with a temperature-density scaling relation of $T\propto n^{-2/3}$, which well approximates the obtained relation in our simulations for the ISM gas below $T<10^{4}\,\rm K$, one gets $p_{\rm sh}=7/12$, steeper than the power of $1/3$ adopted here. ($10^{3}\,\rm H\,cm^{-3}$).
@@ 3.6 Coagulation
Since the velocity dispersion of small grains has complex dependencies with the gas turbulent velocity dispersion at the forcing scale, the magnetic field strength, the grain charge, and the ionisation state of the gas (Yan et al. 2004), we greatly simplify the problem (following Aoyama et al. 2017) by assuming that half ($F=0.5$) the gas mass, which Jeans length is unresolved and with gas density above $0.1\,\rm H\,cm^{-3}$ and temperature below $10^{4}\,\rm K$, has an actually larger gas density of $10^{3}\,\rm H\,cm^{-3}$ with a small grain turbulent velocity dispersion of $0.1\,\rm km\,s^{-1}$ (Yan et al. 2004). We have also performed two simulations with a coagulation model where $F=1$, and where $n$ is sampled by the log-normal PDF shaped by turbulence as for the subgrid model for dust growth by accretion. In this model, we tried two different density cuts, $n_{\rm max,coa}=10^{5}$ or $10^{7}\,\rm H\,cm^{-3}$, and they showed very negligible differences with respect to the fiducial coagulation model for G10LG (they produce respectively $\sim 0.2$ and $3\,\%$ less small grains than the fiducial G10LG simulation at final time).
Figure 3: The dust-to-metal mass ratio (DTM) as a function of time for the fiducial G10LG galaxy. The DTM is further decomposed into the carbonaceous (’C’ in blue) and silicate (’Sil’ in red) grains, and into small $5\,\rm nm$ (’S’ in dotted) and large $0.1\,\mu\rm m$ (’L’ in dashed). The DTM is close to the canonical value of 0.4 in the Milky Way, where the bulk of the mass is composed of large silicate grains.
Figure 4: The dust-to-metal mass ratio (DTM) as a function of time for the different resolutions of the G10LG galaxy using either the fiducial model for dust accretion with turbulence-driven for unresolved densities (solid lines) or the model without (dashed lines). From bottom to top are the simulations with $\Delta x=72$, 36, 18 and $9\,\rm pc$ resolution (the $9\,\rm pc$ resolution simulation is for the dust fiducial model only). The subgrid turbulent model for accretion significantly improves the numerical convergence of the DTM.

### v2

@@ 3.2 Supernova-driven destruction
where $M_{\rm g}$ is the local gas mass, $M_{{\rm d},i,j}$ is the local dust mass, $a_{0.1,j}$ is the dust grain size in units of $0.1\,\rm\mu m$, and $\delta_{{\rm SN},i}$ is the destruction efficiency. Small grains are more efficiently decelerated by drag forces, trapping them near the shock region where thermal sputtering can quickly destroy them (Nozawa et al. 2006). This functional form of the destruction efficiency has been constructed to capture this behaviour with grain size $a$. The mass of shocked gas is provided by the Sedov solution in a medium of homogeneous gas density, i.e. $M_{\rm 100}=6800E_{\rm SN,51}\,\rm M_{\odot}$, where $E_{\rm SN,51}$ is the energy of the SN explosion in units of $10^{51}\,\rm erg$. Multiple SNe were released at once over one time step $\Delta t$ (each individual SN engulfing the gas swept up by the previous explosion), following (see Hou et al. 2017):
@@
with $N_{\rm SN}$ the number of SNe. We stress that this destruction step by the SN blast solution does explicitly destroy the returned ejecta from the stellar production, and destroys only the background dust material, hence, the dust condensation efficiency implicitly takes this term into account. We used a different value of the dust destruction efficiency for carbonaceous and silicate grains with respectively $\delta_{\rm SN,C}=0.10$ and $\delta_{\rm SN,Sil}=0.15$ following the qualitative behaviour of Hu et al. (2019), where silicate grains are approximately destroyed 50 per cent more than carbonaceous grains.
@@ 3.3 Thermal sputtering
Dust is also destroyed by thermal sputtering, the ejection of atoms from grains by the transfer of kinetic energy from gas particles at temperatures high enough to overcome the energy barrier of the binding energy. We adopted the fitting form of Hu et al. (2019) to the thermal sputtering yields $Y_{\rm th}$ calculated by Nozawa et al. (2006), which relates to the thermal sputtering timescale as (recall that $t_{{\rm spu},i,j}=m_{{\rm gr},i,j}/|\dot{m}_{{\rm gr},i,j}|=a_{i,j}/(3|\dot{a}_{i,j}|)$ with $m_{{\rm gr},i,j}$ the mass of a single dust grain):
@@
We note that sputtering yields are approximately 3 times lower at $T\gtrsim 10^{6}\,\rm K$ for carbonaceous grains than for silicate grains (see also Draine & Salpeter 1979; Tielens et al. 1994). Hence, they differ from the widely adopted yields of Tsai & Mathews (1995) giving a unique sputtering timescale based on the intermediate values for silicates and carbonaceous grains of Draine & Salpeter (1979) and Tielens et al. (1994) (see Appendix C). Using cosmological hydrodynamical simulations of galaxy clusters, Gjergo et al. (2018) and Vogelsberger et al. (2019) show that the dust mass content in observed galaxy clusters can be reproduced only for ten times lower sputtering yields than the canonical values. We defer this investigation to future work, but we immediately stress that thermal sputtering is already a sub-dominant destruction mechanism in our set of simulations compared to direct destruction by SNe. We performed a G10LG simulation with sputtering time artificially increased by a factor of 10: there is no noticeable difference in any of the dust properties investigated in section 4.
@@ 3.4 Gas accretion
where $m_{{\rm gr},i,j}=s_{i}a_{j}^{3}4\pi/3$ is the grain mass, $s_{i}$ is the grain material density, $f_{X}$ is the mass fraction of the limiting element in the chemical composition of the grain, $u_{{\rm th},X}=(8k_{\rm B}T/(\pi m_{X}))^{1/2}$ is the gas thermal velocity, $m_{X}=A_{X}m_{\rm H}$ is the atomic mass, and $\rho_{X}$ is the total (gas plus dust phase) mass density of the limiting element. $\pi\bar{a}_{i,j}^{2}$ is the surface of the grain with Coulomb enhancement factor $E_{i}(a_{j})$ due to electrostatic effects caused by ionised gas interacting with charged grains (Weingartner & Draine 1999), which is integrated over the whole distribution of the grain sizes. In our case, where the size distribution was discretised over two bins and assuming a Dirac distribution, we simply got $\bar{a}_{i,j}=E_{i}(a_{j})a_{j}$. For carbonaceous grains, $X=i=\rm C$, however, for silicate grains, the limiting element depends on the grain composition, i.e.
@@
where $a_{{\rm 0.005},j}=a_{j}/5\,\rm nm$, $s_{3,i}=s_{i}/3\,\rm g\,cm^{-3}$, $\alpha$ is the sticking coefficient of gas particles onto dust, and $Z_{X}=\rho_{X}/\rho_{\rm g}$ is the mass fraction of the limiting element. Weingartner & Draine (1999) gave values of $E(a)$ for a range of sizes of carbonaceous and silicate grains for three typical phases of the ISM. To mimic their complex dependencies, we simply adopted a value of $E=1$ for all kind of grains everywhere, except for large carbonaceous and small silicate grains in the cold neutral medium with $T<2\times 10^{4}\,\rm K$ and $n>10\,\rm H\,cm^{-3}$, for which we adopted $E_{\rm C}(0.1\,\mu\rm m)=0$ and $E_{\rm Sil}(5\,\rm nm)=10$.
@@
where $s=\log(n/\bar{n})$, $s_{0}=-\sigma_{s}^{2}/2$ and $\sigma_{s}=\log(1+b^{2}\mathcal{M}^{2})$, with $b$ the compression ratio (we take $b=0.4$ for mixed turbulence which is the same value used in the gravo-turbulent model for SF efficiency) and $\mathcal{M}$ the turbulent Mach number. The effective timescale $t_{\rm acc,eff}$ was obtained by integrating the mass accretion rate over the log-normal PDF for a given set of mean density, temperature, and metallicity ($\bar{n},\bar{T},\bar{Z}$), hence:
@@
up to a maximum density $s_{\rm max}=\log(n_{\rm max}/\bar{n})$, where the maximum density is $n_{\rm max}=10^{4}\,\rm H\,cm^{-3}$ where grains starts to be significantly coated with water ice mantle ($n\geq 10^{3}\,\rm H\,cm^{-3}$, Cuppen & Herbst 2007; Hollenbach et al. 2009) suppressing the accretion of refractory material onto the grain surface. We used the value of the gas density and metallicity of a given cell for $\bar{n}$ and $\bar{Z}$, and we assumed that the gas temperature is always $\bar{T}=100\,\rm K$ for gas that fulfils the conditions to trigger this unresolved log-normal PDF of density (see below) for simplicity. In addition, we assumed that $T=\bar{T}$ and $Z=\bar{Z}$ for the sampled values of $n$ by the log-normal PDF. This led to
@@
where $\rm Erfc$ is the complementary error function. This unresolved effective accretion timescale was estimated only for cells with gas density larger than $0.1\,\rm H\,cm^{-3}$, temperature lower than $10^{4}\,\rm K$, and Jeans length smaller than 4 times the local cell size, assuming a sticking coefficient of $\alpha_{\rm eff}=1/3$, which value is representative of the sticking coefficient obtained at temperatures of $T=100-1000\,\rm K$ (see Fig. 26). Otherwise the accretion time was that obtained for $\bar{n}$ (the local gas density) with the Le Bourlot et al. (2012) sticking coefficient (see Appendix D for a comparison of sticking coefficients).
@@
For carbonaceous grains, we further assumed that carbon in the gas phase is fully molecular (CO) above a gas density of $n>10^{3}\,\rm H\,cm^{-3}$. Indeed, simulations (e.g. Safranek-Shrader et al. 2017; Clark et al. 2019; Hu et al. 2021; Katz et al. 2022) give a transition to full CO gas at $n=2\times 10^{3}-10^{4}\,\rm H\,cm^{-3}$ depending on the cosmic ray ionisation rate. We note that while our value might seem fairly low (or corresponding to a very weak cosmic ray ionisation rate), our subgrid gas distribution neglects the fact that gas collapses further at smaller unresolved scales (higher density) and extends the log-normal PDF with a power-law tail (Vallini et al. 2018).
@@ 3.5 Shattering
Shattering is the process by which grains with sufficiently large velocities collide and fragment into smaller size grains. Therefore, it is a transfer of mass from the large to the small grain population that conserves the total dust mass. The corresponding timescale is obtained by considering the timescale for grain collisions $\tau_{\rm col}=(\sigma_{\rm D}\tilde{\sigma}n_{\rm D})^{-1}$, where $\tilde{\sigma}$ is the grain cross section, $\sigma_{\rm D}$ the dust grain velocity dispersion, and $n_{\rm D}$ the dust number density, which boils down to (Aoyama et al. 2017):
@@
where $\mathcal{D}_{\rm L}$ and $\sigma_{\rm D,L}$ are, respectively, the dust-to-gas ratio and velocity dispersion of the large grains population. Typical velocities required to shatter grains are above a few $\rm km\,s^{-1}$ (Jones et al. 1996), which limit the effect of shattering to the diffuse phase of the ISM. Yan et al. (2004) provided grain-gas relative velocities for various phases of the ISM and showed that large grains ($a_{\rm L}=0.1\,\rm\mu m$) reach the turbulent velocity of the forcing scale, i.e. $\sigma_{\rm D,L}\simeq 10\,\rm km\,s^{-1}$ in the warm ionised medium, $\simeq 1\,\rm km\,s^{-1}$ for the warm or cold neutral medium, and $\simeq 0.1-1\,\rm km\,s^{-1}$ for molecular clouds. We followed Granato et al. (2021) and adopted the following functional form:
@@
with $p_{\rm sh}=1$ for $n<1\,\rm H\,cm^{-3}$ and $p_{\rm sh}=1/3$ for $1<n<10^{3}\,\rm H\,cm^{-3}$, which is equivalent to having a dispersion velocity varying from $10\,\rm km\,s^{-1}$ to $1\,\rm km\,s^{-1}$ from the warm ionised phase ($1\,\rm H\,cm^{-3}$) to the molecular cloud density [^5] We note that considering the gas-drag timescale and a Kolmogorov turbulence, the velocity dispersion of grains follow (Ormel et al. 2009; Hirashita & Aoyama 2019): $\sigma_{\rm gr}=1.2\mathcal{M}^{3/2}a_{0.1}^{1/2}s_{3}^{1/2}T_{4}^{1/4}n_{1}^{-1/4}\,\rm km\,s^{-1}$. Hence, with a temperature-density scaling relation of $T\propto n^{-2/3}$, which well approximates the obtained relation in our simulations for the ISM gas below $T<10^{4}\,\rm K$, one gets $p_{\rm sh}=7/12$, steeper than the power of $1/3$ adopted here. ($10^{3}\,\rm H\,cm^{-3}$).
@@ 3.6 Coagulation
Since the velocity dispersion of small grains has complex dependencies with the gas turbulent velocity dispersion at the forcing scale, the magnetic field strength, the grain charge, and the ionisation state of the gas (Yan et al. 2004), we greatly simplified the problem (following Aoyama et al. 2017) by assuming that half ($F=0.5$) the gas mass, which Jeans length is unresolved and with gas density above $0.1\,\rm H\,cm^{-3}$ and temperature below $10^{4}\,\rm K$, has an actually larger gas density of $10^{3}\,\rm H\,cm^{-3}$ with a small grain turbulent velocity dispersion of $0.1\,\rm km\,s^{-1}$ (Yan et al. 2004). We also performed two simulations with a coagulation model where $F=1$ and where $n$ is sampled by the log-normal PDF shaped by turbulence as for the subgrid model for dust growth by accretion. In this model, we tried two different density cuts, $n_{\rm max,coa}=10^{5}$ or $10^{7}\,\rm H\,cm^{-3}$, and their results show very negligible differences with respect to the fiducial coagulation model for G10LG (they produce respectively $\sim 0.2$ and $3\,\%$ less small grains than the fiducial G10LG simulation at final time).
Figure 3: Dust-to-metal mass ratio (DTM) as a function of time for the fiducial G10LG galaxy. The DTM is further decomposed into the carbonaceous (’C’ in blue) and silicate (’Sil’ in red) grains, and into small $5\,\rm nm$ (’S’ in dotted) and large $0.1\,\mu\rm m$ (’L’ in dashed). The DTM is close to the canonical value of 0.4 in the Milky Way, where the bulk of the mass is composed of large silicate grains.
Figure 4: Dust-to-metal mass ratio (DTM) as a function of time for the different resolutions of the G10LG galaxy using either the fiducial model for dust accretion with turbulence-driven for unresolved densities (solid lines) or the model without (dashed lines). From bottom to top are the simulations with $\Delta x=72$, 36, 18, and $9\,\rm pc$ resolution (the $9\,\rm pc$ resolution simulation is for the dust fiducial model only). The subgrid turbulent model for accretion significantly improves the numerical convergence of the DTM.